# The optimum mix of storage and backup in a highly renewable, highly reliable European electricity grid

### v1

@@ Abstract
Computationally expensive models have other limitations. Previous studies have looked at the effects of carbon-intensity and individual technology costs on the future grid [12, 13], but understanding how these factors combine to affect the overall system cost requires searching a high-dimensional space. For complex models, this can be prohibitively expensive.
@@
In this paper, we give results from a simplified model for the European power sector able to resolve some of these more difficult questions. Inspired by the work of Shaner et. al. [14] and Tong et. al. [15], we use the MERRA-2 reanalysis datasets to estimate solar and onshore-wind generation potential in Europe over a 43-year period from 1980–2022. We model Europe as a single node, averaging supply and demand across the continent and assuming perfect, lossless transmission. We then simulate this idealised grid with hourly resolution from 1980–2022, varying the installed capacities of solar, onshore wind, storage and dispatchable backup as exogenous inputs. With this single-node, four-technology grid model, we can investigate how three parameters—the permitted carbon intensity of the system, the cost of electricity storage, and the grid reliability target—interact to affect the overall system cost. These constraints are highly interdependent, and must be studied together in order to fully capture their impact on the system cost. To estimate the costs of storage and dispatchable backup, we assume lithium-ion batteries, due to their widespread adoption, and natural gas, due to its low capital cost. The model is described in further detail in the Methods section.
.2 Limitations of Our Model
This model allows new questions to be answered, but it has limitations. Because we do not consider transmission and only model four technologies, the total system costs reported here are not realistic estimates for true European system costs. Specifically, including transmission costs would increase costs while including alternative and existing sources of generation would decrease the overall system cost.
Furthermore, while the model uses 43 years of supply data, a lack of data availability means that we only include electricity demand data for the year 2022. The consequences of this are discussed in detail in the Supplementary Information, including an analysis of potential correlations between renewable-energy generation and electricity demand that are not captured by our analysis.
Finally, there is an extensive discussion in the literature surrounding the validity of reanalysis datasets like MERRA-2 for renewable generation studies [6, 7, 16, 17]. Staffell and Pfenninger found discrepancies between wind and solar generation estimates from MERRA-2 and real-world generation data for European countries [6, 7]. To validate the conclusions of this paper, we have performed our analysis using both uncorrected MERRA-2 data as well as renormalised datasets for wind and solar published by Staffell and Pfenninger on the renewables.ninja platform [6, 7]. The countries covered in our original analysis and the countries covered in our validation analysis with the renewables.ninja dataset are shown in Fig. 1. Full details of the differences between the datasets are included in Appendix A, and a direct comparison of the results between the two datasets is included in the Supplementary Information. The results and conclusions presented in this paper were found to be robust to the choice of dataset used.
I Required Overbuild & Storage to Meet Reliability Targets
Figure 2 shows the daily and seasonal variation in renewable generation over the 43-year period. Solar output is more predictable than wind output, but experiences much larger variation over the course of a day (Fig. 2a) and over the course of a year (Fig. 2b). Figure 2b also highlights the complementarity between wind and solar when averaged across the European continent, since solar is strongly peaked in summer and wind is peaked in winter. This supports previous studies, which have found complementarity between seasonal wind and solar generation in Europe [18, 19].
Fig. 2 also indicates that, without any overbuild, solar and wind rarely meet demand at any hour of the year. This is unsurprising, because current European capacity factors are between 10%–20% for solar and 10%–30% for onshore wind [18]. In Fig. 3a, we quantify this by plotting the obtained grid reliability over the 43-year period for different renewable overbuild factors as a function of the percentage of wind power in the generation mix. For overbuild factors of $\times$1, meaning that the combined nameplate capacity of wind and solar equals peak demand, demand is almost never met for any mix of wind and solar. Even when the installed wind and solar capacity is 4860 GW—ten times the peak electricity demand—the resulting grid reliability of 99.94% still falls short of modern 99.97% reliability targets. Diminishing returns are also observed at high overbuild factors: increasing the overbuild factor from $\times$7 to $\times$10 of peak demand only boosts grid reliability 1.4% from 98.54% to 99.94%. [^2] It should be noted that the interpretation of the renewable overbuild factors presented here is different from the interpretation in Shaner et. al. [14] and Tong et. al. [15]. In these papers, $\times$1 overbuild is defined as ‘average renewable generation = average demand’, whereas in our analysis $\times$1 overbuild is defined as ‘installed renewable capacity = peak demand’, which we believe is a more useful metric in helping to size the system.
Adding energy storage into the mix, operated according to the rules specified in the Methods section, rapidly reduces the required overbuild factors to reach target levels of grid reliability. Fig. 3b plots the renewable overbuild factor required to meet different reliability targets as a function of the storage capacity in the system. For a 99.97% reliability target, 1458 GWh of storage (equivalent to 3 hours at peak demand) is sufficient to reduce the required overbuild factor from $\times$10.6 to $\times$7. However, once storage capacity extends beyond a few hours of peak demand, it has less of an effect on the required overbuild factor. A full 24 hours of storage (11,664 GWh) only reduces the required overbuild factor for 99.97% reliability to $\times$5.
Installed battery storage capacity in Europe currently sits at about 17.2 GWh [20], which corresponds to two minutes of peak demand. However, this number has doubled every year since 2020, and significant additional storage capacity is planned. Our results suggest that the first 1000 GWh of storage yield the greatest benefits to the system, with diminishing returns after this. Interestingly, pumped hydropower storage already provides $\sim$2000 GWh in Europe—although 38% of this is located in Norway alone and a further 17% is in Switzerland [21]. Geopolitical issues notwithstanding, the large-scale storage provided by pumped hydropower will be a crucial component in a future, highly renewable European electricity grid.
Solar 790 [22] 10 [23] 25 [22]
Wind 1540 [22] 40 [23] 25 [22]
Storage 200 (\$/kWh) [24] 10 (\$/kWh-year) [24] 15 [24]
Natural Gas 1000 [22] 20 [25] 30 [26]
Table 1: Cost estimates for each technology used in the analysis. Sources are IEA [22], IRENA [23], EIA [25] and NREL [24, 26]. Variable Operations & Maintenance (O&M) costs were assumed negligible for all technologies, and fuel costs for natural gas were neglected because its maximum utilisation was just 2% of all electricity generation in the study. All cost data in \$$\$_{2024}$. Fig. 5 shows a sensitivity analysis to storage costs. The conversion to investment cost is described in the Methods section.
II Cost-optimised scenarios
Previous studies examining the interplay of storage and dispatchable backup in Europe [27] and the United States [28] only dispatched backup generation when storage levels fell to zero. However, it is plausible that deploying dispatchable backup before the storage levels reach zero could improve grid reliability or reduce the backup capacity required. To allow for this possibility, we adopt a threshold model for the use of natural gas backup (described in the Methods section), wherein natural gas is dispatched once storage levels drop below a pre-specified threshold.
For a given grid configuration, the simulation returns the average grid reliability over the 43-year period and the percentage of electricity that was produced by natural gas.
We combine this with a cost model for the system (see Table 1 and the Methods section), and find the minimum-cost blend given constraints on grid reliability and natural-gas usage (a proxy for carbon intensity if not paired with carbon capture and storage). We do not include transmission in our cost estimates, since we do not model transmission, but achieving the perfect levels of interconnectivity assumed here would require major upgrades to the European transmission grid.
II.1 System costs vs. grid reliability
Figure 4 shows how the annual system cost varies as a function of the reliability target for 3 different constraints on natural gas usage, with the different colours in Fig. 4a representing different levels of permitted natural gas generation. Figs. 4b–e contain further details about the cost-optimised results for the four circled cases in 4a.
The red line in Figure 4a indicates that, in a European energy system containing only wind, solar and storage, the cost difference between a 99% reliable grid and a 100% reliable grid is about \$200 billion annually—or 36% of the entire system cost. As soon as natural gas backup is added into the mix, not only are major cost savings achieved, but the cost savings associated with relaxed grid reliability targets are also reduced. If 1% of total electricity generation is permitted to come from natural gas, then the optimal cost of a 99.97% reliable grid falls from \$491 billion annually to \$335 billion annually. Furthermore, with 1% of electricity from natural gas, this 99.97% reliable grid can be achieved at relatively little additional cost compared with a 99% reliable grid (just \$13 billion extra per year). These conclusions are consistent with the results in [12], which found that the costs of decarbonising the power sector in the absence of firm resources grow rapidly for the last few percent of generation.
Figure 5: a) Optimised annual system cost versus storage capital cost for a 99.97% reliable grid. b) is 0% natural gas at storage cost of \$10 per kWh; c) is 0% natural gas at storage cost of \$400 per kWh; c) is 1% natural gas at storage cost of \$10 per kWh; d) is 1% natural gas at storage cost of \$400 per kWh. The subplots display the same information as in Fig. 4.
From the cost breakdowns in Figure 4di and Figure 4ei, it is clear that the \$13 billion cost of increasing the reliability from 99% to 99.97% is the result of a doubling of natural gas capacity (not usage)—from 121 GW to 218 GW—while the rest of the infrastructure required in the cost-optimised case remains largely the same. When natural gas is allowed into the system, the installed natural gas capacity becomes the primary lever for increasing grid reliability.
The other difference between Figure 4di and Figure 4ei is an increase in the optimum threshold level determined by the simulation—from 0% to 5%. To reach very high reliability levels, the backup capacity must start being deployed before the storage levels fall to zero. We would expect the optimum threshold level to rise further under stricter reliability targets or weaker carbon constraints.
It is also interesting to note that wind significantly dominates the overall cost in all scenarios, and that all scenarios rely on large renewable overbuild factors ($\times$7 for Fig. 4b, $\times$5.8 for Fig. 4c, $\times$5.2 for Fig. 4d and $\times$5 for Fig. 4e, defined as nameplate capacity over peak demand), even with storage and natural gas backup in the system. At the end of 2023, total installed wind capacity in Europe was 272 GW, with annual additions of 20 GW in 2022 and 18 GW in 2023 [29, 30]. At a build-rate of 20 GW per year, total installed capacity would be 812 GW in 2050, which is at least a factor of two lower than the wind capacity required to achieve a reliable grid in all our scenarios.
European installed solar capacity in Europe in 2023 was similar (263 GW), but annual additions in 2023 were much higher than for wind at 56 GW [31]. At this build-rate, there would easily be enough solar capacity to cover all of our scenarios in Figs. 4 and 5. This highlights a fundamental tension in the deployment of renewable energy in Europe: solar projects are more profitable than wind at an individual level (hence the rapid roll-out), but the overall system costs are reduced when wind is prioritised.
II.2 System costs vs. storage capital costs
Figure 5 shows instead how the optimised system cost for a 99.97% reliable grid varies as a function of the storage capital cost. In the results shown in Fig. 4, we assumed \$200 per kWh(following [24]), but there is a large uncertainty in this estimate for 2050 battery storage costs. We explore here how the total system cost is affected by the value chosen for the storage capital cost. Again, we include details for four specific example cases in Figs. 5b-e.
These results show that the cost reductions arising from the inclusion of natural gas as a firm and flexible backup generator significantly outweigh the cost reductions from any decrease in storage cost. For a 0% natural gas grid, reducing the capital cost of storage from \$400 per kWh down to \$10 per kWh results in annual system cost savings of \$93 billion (\$515 billion to \$422 billion). Allowing 1% of electricity to come from natural gas leads to annual system cost savings of \$168 billion (\$515 billion to \$347 billion), even when storage costs are kept fixed at \$400 per kWh. Indeed the cost savings that result from a further 1% increase in allowed natural gas generation are comparable to the cost savings from reducing the storage capital costs from \$400 per kWh to \$10 /kWh again.
The overall system costs become less sensitive to storage capital costs when natural gas is allowed in the system. Without gas, increasing storage capital costs from \$10 per kWh to \$400 per kWh increases system costs by 22%, compared to a 12% increase with 1% gas in the system.
It is worth noting that significant storage capacity is still included in the cost-optimised scenario at \$400 per kWh, even when natural gas backup is permitted. This suggests that some storage is crucial to a low-cost system and supports the conclusions of [27], which found that adding small amounts of storage to the grid dramatically reduces the energy that must be covered by dispatchable backup resources.
The patterns of outages under varying storage costs are similar to those seen in Fig. 4. For cases 5d and 5e with natural gas permitted, there are no outages over the entire 43-year period in which less than 70% of demand is met. Without natural gas, however, cases 5b and 5c show that hours where less than 70% of demand is met do occur. Furthermore, Fig. 5biii shows one outage in which supply falls short of demand for over 24 hours continuously. When natural gas is added to the system, no outage lasts longer than 10 hours.
III Conclusions
Using a simplified model, we have explored the trade-offs required to achieve a highly reliable, low-carbon electricity grid in Europe. Our findings indicate that a reliable renewables-dominated European grid will require high levels of overbuild, with installed peak generating capacity greater than four times the peak demand. Even so, we find that cost-minimised systems require both storage and dispatchable backup generation to achieve reliability targets. A large quantity of one cannot fully compensate for the complete absence of the other without incurring extreme penalties in system costs.
Our study suggests a number of important policy findings. First, grid optimisation studies that force a zero-emissions target are likely to produce conclusions in terms of technology mix, cost, and feasibility that deviate significantly from the true social optimum, as disallowing very small emissions has a major impact on cost and social welfare. As such, finding ways to preserve some legacy generators as backup capacity may be valuable if a renewables-dominated grid is to be pursued.
Finally, it is important to note that the cost estimates presented in this study are illustrative. While they provide insight into cost differentials across different technology scenarios, they do not represent total system costs. A comprehensive estimate would need to account for sunk infrastructure costs, transmission expansion, and a broader range of generation technologies. Technology-rich studies, albeit with shorter temporal coverage and poorer reliability insight, can be found elsewhere in the literature.
IV Methods
@@
The simulations in this analysis rely on hourly time series data for solar generation and wind generation over a 43-year period (1980–2022), which were generated using raw weather data from the MERRA-2 reanalysis dataset [32, 33]. Hourly electricity demand data for the year 2022 was collated from the ENTSO-E Transparency Platform [34].
@@
IV.1 Calculating solar generating potential
In order to calculate solar generation at every grid point, we used the incident short-wave radiation at ground level to compute direct illumination (variable name: SWGDN) and the incident radiation at the top of the atmosphere to compute diffusion illumination (variable name: SWTDN) from the Radiation Diagnostics dataset [32]. We use the 2-metre ambient air temperature (variable name: T2M) from the Single-Level Diagnostics dataset [33] to adjust cell efficiency.  [^3] Note that SWGDN from the Radiation Diagnostics dataset differs from the SWGDN values in the Land Surface Forcings dataset [35] as the incident radiation is averaged over the total area of each grid cell in the Radiation Diagnostics dataset, while it is only averaged over the land area in each grid cell in the Land Surface Forcings dataset.
To convert these variables to a solar output, we assumed non-tracking solar panels and followed a similar method to Pfenninger and Staffell [7] and Sandia National Laboratory [36]. First, we calculated the in-plane solar irradiation of the panel
$I_{tot}=I_{dir,in\text{---}plane}+I_{dif,in\text{---}plane}+I_{refl,in\text{---}plane},$ (1)
@@ IV.1 Calculating solar generating potential
$I_{refl,in\text{---}plane}=\textit{GHI}\times\frac{a(1-\cos{t})}{2},$ (2)
@@
The diffuse fraction of the total GHI was estimated using the logistic model proposed in [37] as
$\textit{DHI}=\frac{\textit{GHI}}{1+e^{-5.0033+8.6025k_{t}}},$ (3)
where $k_{t}$ is the clearness index, which we estimate as the ratio of the ground-level radiation (SWGDN) to extraterrestrial radiation (SWTDN), and DHI is the Diffuse Horizontal Irradiation—the diffuse irradiation measured normal to the Earth’s surface. This was converted to an in-plane value via
$I_{dif,in\text{---}plane}=\textit{DHI}\times\frac{1+\cos{(t)}}{2}.$ (4)
We calculated the direct component of the solar radiation using
@@
The in-plane component of the direct radiation, $I_{dir,in\text{---}plane}$, is then given by
$I_{dir,in\text{---}plane}=\textit{DNI}\cos{(\textit{AOI})},$ (6)
@@
We accounted for the temperature dependence of solar panel efficiency following Huld, [38] assuming crystalline silicon solar cells and a relationship between panel temperature and ambient air temperature of $T_{panel}=T_{amb}+0.035\times I_{in\text{---}plane}$, with ambient air temperature from the MERRA-2 T2M variable.
We assumed a standard efficiency (at $I_{in\text{---}plane}$ = 1000 $W/m^{2}$ and $T_{panel}$ = 25 °C) of 21% [39] for all solar panels. The power output $P$ of the panel in $W/m^{2}$ can then be calculated as
$P=0.21\eta_{rel}I_{in\text{---}plane},$ (8)
@@
IV.2 Calculating wind generating potential
We calculated the wind power generated over the 43-year period from a typical modern wind turbine placed in each grid cell. We assumed a turbine rotor diameter of 110 m, a hub-height of 100 m, a maximum rated power of 4.1 MW and cut-in and cut-out speeds of 3 m/s and 25 m/s respectively.
@@ IV.2 Calculating wind generating potential
To estimate $v$ at the hub height of 100 m, we used the 10-metre wind speeds (variable names V10M and U10M) and the 50-metre wind speeds (V50M and U50M) from the MERRA-2 Single-Level Diagnostics dataset [33]. We used a power law relationship of the form
@@
to extrapolate these wind speeds to the hub height of 100 m [41]. In Eq. 10, $v_{h_{1}}$ and $v_{h_{2}}$ are the wind speeds at heights $h_{1}$ and $h_{2}$, and $a$ is a constant that depends on site-specific geography. We estimated the atmospheric density $\rho_{air}$ at ground level from the the 2-metre air temperature (variable name T2M) and the surface pressure (variable name PS), and extrapolated to the hub height of 100 m using the barometric equation [42]. We combined Eq. 9 with the constraints on turbine operation (cut-in speed, cut-out speed, and maximum output) to create a piece-wise model for wind output at every grid square:
@@
We used these expressions to obtain a ‘wind power per turbine’ and a ‘solar power per unit area’ at every point on the MERRA-2 grid. We then averaged over all 2044 points in the grid to obtain a Europe-wide ‘wind power per turbine’ and ‘solar power per unit area’ for every hour in the dataset.
IV.3 Estimating electricity demand
The European Network of Transmission System Operators for Electricity (ENTSO-E), which operates the synchronous grid in Europe, requires its members to publish hourly electricity demand data every year, and makes this information publicly available [34]. However, not all countries connected to the synchronous European grid (CESA) are members of ENTSO-E, and the UK, Turkey and the North African states in particular do not publish hourly demand data on the ENTSO-E platform. Data for the UK are available elsewhere [43], but Turkey and the North African states were excluded from our analysis due to data being unavailable. Some countries, such as Ukraine and Albania, use the ENTSO-E platform but do not reliably publish data. We only included in our analysis the 36 countries that provided demand data for more than 8700 hours of the year (out of 8760) in 2022—see Appendix B. The borders of these countries define the locations of the MERRA-2 grid points used to calculate solar and wind generating potential.
We augmented demand data for missing hourly demand values with linear interpolation. To model demand over the full 43-year period of the MERRA-2 dataset, we duplicated the 2022 demand data for each year 1980–2022, as demand data is not available over the full period covered by MERRA-2. For a detailed discussion of this point, and further analysis of the consequences of using a single year of demand data, see the Supplementary Information.
IV.4 Storage & Backup
In order to analyse the effects of adding storage and dispatchable backup to a highly renewable electricity grid, we had to choose technologies to capture these characteristics. For storage, we chose lithium-ion batteries, since this is currently the lowest cost battery technology and has been widely studied in the context of future grid-scale electricity storage [44, 45]. We followed the cost estimates in [24] for a 4-hour Li-ion storage system, and assumed that maintenance covered under the fixed annual O&M costs ensures no degradation of the battery over its life cycle [24]. For simplicity, we have assumed perfect round-trip efficiency and no self-discharge for the batteries. We would not expect any significant changes to our results if these terms were included since, in our threshold-based model, storage inefficiencies would lead to natural gas backup being deployed slightly earlier. This would increase fossil fuel emissions for each scenario slightly, but would likely not affect the system cost or the reliability of the grid.

### v2

@@ Abstract
keywords:
Long-term grid reliability , Energy system modelling , Optimising renewable generation mix , Net Zero in Europe , Decarbonisation trade-offs
{highlights}
First multi-decadal simulation of a renewable European electricity grid
Both storage and dispatchable backup are crucial for maintaining grid reliability
Dispatchable backup can enable large cost-savings at low utilisation rates
Cost-optimised scenarios are dominated by wind rather than solar
1 Introduction
@@
Computationally expensive models have other limitations. Previous studies have looked at the effects of carbon-intensity constraints and individual technology costs on the future grid [12, 13], but understanding how these factors combine to affect the overall system cost requires searching a high-dimensional space. For complex models, this can be prohibitively expensive.
@@
In this paper, we give results from a reduced model for the European power sector capable of resolving some of these more difficult questions. Inspired by the work of Shaner et. al. [14] and Tong et. al. [15], we use the MERRA-2 reanalysis dataset to estimate solar and onshore-wind generation potential in Europe over a 43-year period from 1980–2022. We model Europe as a single node, averaging supply and demand across the entire continent and assuming perfect, lossless transmission. We then simulate this idealised grid with hourly resolution from 1980–2022, varying the installed capacities of solar, onshore wind, storage and dispatchable backup as exogenous inputs. With this single-node, four-technology grid model, we can investigate how three parameters—the permitted carbon intensity of the system, the cost of electricity storage, and the grid reliability target—interact to affect the overall system cost. These constraints are highly interdependent, and must be studied together in order to fully capture their impact on the system cost.
2 Methods I - Data
@@
The simulations in this analysis rely on hourly time series data for solar generation and wind generation over a 43-year period (1980–2022). These time-series were generated using raw weather data from the MERRA-2 reanalysis dataset [16, 17]. Hourly electricity demand data for the year 2022 was collated from the ENTSO-E Transparency Platform [18].
@@
2.1 Calculating solar generating potential
In order to calculate solar generation at every grid point, we used the incident short-wave radiation at ground level (variable name: SWGDN) to compute direct illumination and the incident radiation at the top of the atmosphere to compute diffusion illumination (variable name: SWTDN). Both variables are contained within the MERRA-2 Radiation Diagnostics dataset [16]. We used the 2-metre ambient air temperature (variable name: T2M) from the Single-Level Diagnostics dataset [17] to account for the temperature-dependence of cell efficiency.  [^2] Note that SWGDN from the Radiation Diagnostics dataset differs from the SWGDN values in the Land Surface Forcings dataset [19] as the incident radiation is averaged over the total area of each grid cell in the Radiation Diagnostics dataset, while it is only averaged over the land area in each grid cell in the Land Surface Forcings dataset.
To convert these variables into a solar output, we assumed non-tracking solar panels and followed a similar method to Pfenninger and Staffell [7], and Sandia National Laboratory [20]. First, we calculated the in-plane solar irradiation of the panel
$I_{tot}=I_{dir,in\text{-}plane}+I_{dif,in\text{-}plane}+I_{refl,in\text{-}plane},$ (1)
@@ 2.1 Calculating solar generating potential
$I_{refl,in\text{-}plane}=\textit{GHI}\times\frac{a(1-\cos{t})}{2},$ (2)
@@
In order to calculate $I_{dif,in\text{-}plane}$ in Eq. 1, we first estimated the diffuse fraction of the total GHI using the logistic model proposed in [21]:
$\textit{DHI}=\frac{\textit{GHI}}{1+e^{-5.0033+8.6025k_{t}}}.$ (3)
Here, $k_{t}$ is the clearness index, which we estimated as the ratio of the ground-level radiation (SWGDN) to extraterrestrial radiation (SWTDN), and DHI is the Diffuse Horizontal Irradiation—the diffuse irradiation measured normal to the Earth’s surface. This was converted to an in-plane value via
$I_{dif,in\text{-}plane}=\textit{DHI}\times\frac{1+\cos{(t)}}{2}.$ (4)
To obtain $I_{dir,in\text{-}plane}$ in Eq. 1, we calculated the direct component of the solar radiation using
@@
The in-plane component of the direct radiation, $I_{dir,in\text{-}plane}$, is then given by
$I_{dir,in\text{-}plane}=\textit{DNI}\cos{(\textit{AOI})},$ (6)
@@
We accounted for the temperature dependence of solar panel efficiency following Huld [22], assuming crystalline silicon solar cells and a relationship between panel temperature and ambient air temperature of $T_{panel}=T_{amb}+0.035\times I_{in\text{-}plane}$, with ambient air temperature from the MERRA-2 T2M variable.
We assumed a standard efficiency (at $I_{in\text{-}plane}$ = 1000 $W/m^{2}$ and $T_{panel}$ = 25 °C) of 21% [23] for all solar panels. The power output $P$ of the panel in $W/m^{2}$ can then be calculated as
$P=0.21\eta_{rel}I_{in\text{-}plane},$ (8)
@@
2.2 Calculating wind generating potential
For wind generating potential, we calculated the wind power that would be generated over the 43-year period by a typical modern wind turbine placed in each grid cell. We assumed a turbine rotor diameter of 110 m, a hub-height of 100 m, a maximum rated power of 4.1 MW and cut-in and cut-out speeds of 3 m/s and 25 m/s respectively.
@@ 2.2 Calculating wind generating potential
To estimate $v$ at the hub height of 100 m, we used the 10-metre wind speeds (variable names V10M and U10M) and the 50-metre wind speeds (V50M and U50M) from the MERRA-2 Single-Level Diagnostics dataset [17]. A power law relationship of the form
@@
was then used to extrapolate these wind speeds to the hub height of 100 m [25]. In Eq. 10, $v_{h_{1}}$ and $v_{h_{2}}$ are the wind speeds at heights $h_{1}$ and $h_{2}$, and $a$ is a constant that depends on site-specific geography. We estimated the atmospheric density $\rho_{air}$ at ground level from the the 2-metre air temperature (variable name T2M) and the surface pressure (variable name PS), and extrapolated to the hub height of 100 m using the barometric equation [26]. We combined Eq. 9 with the constraints on turbine operation (cut-in speed, cut-out speed, and maximum output) to create a piece-wise model for wind output at every grid square:
@@
Eqs. 8 and 11 were used to obtain a ‘solar power per unit area’ and a ‘wind power per turbine’ at every point on the MERRA-2 grid. We then averaged over all 2044 points in the grid to obtain a Europe-wide ‘wind power per turbine’ and ‘solar power per unit area’ for every hour in the dataset.
2.3 Estimating electricity demand
The European Network of Transmission System Operators for Electricity (ENTSO-E), which operates the synchronous grid in Europe, requires its members to publish hourly electricity demand data every year, and makes this information publicly available [18]. However, not all countries connected to the synchronous European grid (CESA) are members of ENTSO-E, and the UK, Turkey and the North African states in particular do not publish hourly demand data on the ENTSO-E platform. Data for the UK are available elsewhere [27], but Turkey and the North African states were excluded from our analysis due to data being unavailable. Some countries, such as Ukraine and Albania, use the ENTSO-E platform but do not reliably publish data. We only included in our analysis the 36 countries that provided demand data for more than 8700 hours of the year (out of 8760) in 2022—see B. The borders of these countries define the locations of the MERRA-2 grid points used to calculate solar and wind generating potential.
For the remaining hours with missing demand data, we estimated these values using linear interpolation. To model demand over the full 43-year period of the MERRA-2 dataset, we duplicated the 2022 demand data for each year 1980–2022, as demand data is not available over the full period covered by MERRA-2. For a detailed discussion of this point, and further analysis of the consequences of using a single year of demand data, see the Supplementary Information.
2.4 Storage & Backup
In order to analyse the effects of adding storage and dispatchable backup to a highly renewable electricity grid, we had to choose technologies to capture these characteristics. For storage, we chose lithium-ion batteries, since this is currently the lowest cost battery technology and has been widely studied in the context of future grid-scale electricity storage [28, 29]. We followed the cost estimates in [30] for a 4-hour Li-ion storage system, and assumed that maintenance covered under the fixed annual O&M costs ensures no degradation of the battery over its life cycle [30]. For simplicity, we have assumed perfect round-trip efficiency and no self-discharge for the batteries. We would not expect any significant changes to our results if these terms were included since, in our threshold-based model (see Section 3), storage inefficiencies would lead to natural gas backup being deployed slightly earlier. This would increase fossil fuel emissions for each scenario slightly, but would likely not affect the system cost or the reliability of the grid.